Suppose a special type of small data processing firm is so specialized that some have difficulty making a profit in their first year operation. The pdf that characterizes proportion Y that makes the profit is given by:

f(y)= Ky(1-y)2 0< y < 1
(a): What is the value of “K” that renders the above a valid density function?
(b): Find the probability that at most 40% of the firms make a profit in the first year.
1
Expert's answer
2020-06-03T19:22:39-0400

We have the pdf that characterizes proportion Y that makes the profit is given by:

f(y)= Ky(1-y)2 , 0< y < 1


(a) Now, f(y) will be a valid density function if,


"\int_0^1f(y)dy=1"


"\implies\int_0^1Ky(1-y)^2dy=1"


"\implies K\int_0^1(y-2y^2+y^3)dy=1"

"\implies K[\int_0^1ydy-2\int_0^1y^2dy+\int_0^1y^3dy]=1"


"\implies K[\frac{1}{2}y^2-\frac{2}{3}y^3+\frac{1}{4}y^4]_0^1=1"


"\implies K[\frac{1}{2}-\frac{2}{3}+\frac{1}{4}]=1"


"\implies K.\frac{1}{12}=1"


"\implies K = 12"


Answer: The value of K = 12 will render the given function a valid density function.


(b) The probability that at most 40% of the firms make a profit in the first year


P(Y "\leq" 40%)


= P(Y "\leq" 0.4)


= "\int_0^{0.4}f(y)dy"


= "12\int_0^{0.4}y(1-y)^2dy"


= "12[\int_0^{0.4}ydy-2\int_0^{0.4}y^2dy+\int_0^{0.4}y^3dy]"


= "12[\frac{1}{2}y^2-\frac{2}{3}y^3+\frac{1}{4}y^4]_0^{0.4}"


= "12[\frac{1}{2}\times{0.4}^2-\frac{2}{3}\times{0.4}^3+\frac{1}{4}\times{0.4}^4]"


= 0.5248


Answer: The probability that at most 40% of the firms make a profit in the first year is 0.5248.